OpenStudy (anonymous):

The red bus company charges $150 plus $72 per hour to rent a bus. The blue bus company charges $240 plus $54 per hour. For what number of hours would the companies charge the same? What would the charge be for that number of hours?

OpenStudy (anonymous):

make to equations and set them equal to each other:\[\text{let h = # of hours}\]\[150+72h=240+54h\]solve for h\[150+18h=240\]\[18h=90\]\[\large{h=5}\]now plug in 5 into either one of the two equations:\[150+72(5)\]\[150+360\]\[\large{510}\] so... \[\large\text{At 5 hours both companies will be charging an equal amount of}\]\[\large\text{510 dollars}\] hope this helps! :)

OpenStudy (anonymous):

Thanks so much!! can you maybe help me with this one too? Suppose it is 40 degrees inside your refrigerator and 360 degrees inside your oven when the power goes out. The refrigerator temperature starts rising at the rate of 5 degrees per hour. the oven temperature starts dropping at the rate of 45 degrees per hour. How long will it take for the temperatures to be the same? What will the temperature be then?

OpenStudy (anonymous):

so sorry....i had logged out...i dont know if this is still of help but ill do the question anyway :) \[\text{Let \(h\) = the # of hours since the power went out}\]now lets make two equations and set them equal to each other:\[40+5h=360-45h\]now just solve for \(h\):\[40+50h=360\]\[50h=320\]\[h=6.4\]now just plug it into either one of the two equations:\[40+5(6.4)\]\[40+32\]\[72\] so...\[\large\text{The temperatures will be the same at 72 degrees after 6.4 hrs}\]